# Exact Quench Dynamics from Thermal Pure Quantum States

### v1

@@ D.3 The BCS representation of the crosscap state
$H_{\text{parent}}=\sum_{k>0}\Psi_{k}^{\dagger}\mathbf{H}_{k}\Psi_{k},$ (S118)
@@
$\mathbf{H}_{k}=\begin{pmatrix}\epsilon_{k}&\Delta_{k}\\ \Delta_{k}^{*}&-\epsilon_{k}\end{pmatrix}=\begin{pmatrix}0&i\sin(kN)\\ -i\sin(kN)&0\end{pmatrix}\,$. (S119)
Here, we’ve used the specific values for our parent Hamiltonian: the diagonal ”energy” term is $\epsilon_{k}=0$, and the off-diagonal ”pairing” term is $\Delta_{k}=i\sin(kN)$.
@@
We find the coefficients by substituting the transformation into the Hamiltonian and demanding that all off-diagonal terms like $b_{k}^{\dagger}b_{-k}^{\dagger}$ and $b_{k}b_{-k}$ vanish. This leads to a standard set of equations that relate the coefficients to the Hamiltonian parameters $\epsilon_{k}$ and $\Delta_{k}$.
@@
$\mathcal{E}_{k}=\sqrt{\epsilon_{k}^{2}+|\Delta_{k}|^{2}}$ (S121)
@@
$|u_{k}|^{2}=\frac{1}{2}\left(1+\frac{\epsilon_{k}}{\mathcal{E}_{k}}\right)\quad\text{and}\quad|v_{k}|^{2}=\frac{1}{2}\left(1-\frac{\epsilon_{k}}{\mathcal{E}_{k}}\right)\,$, (S122)
@@
$\frac{v_{k}}{u_{k}}=\frac{\Delta_{k}}{\epsilon_{k}+\mathcal{E}_{k}}\,.$ (S123)
Now we substitute the specific parameters for the crosscap state’s parent Hamiltonian: $\epsilon_{k}=0$ and $\Delta_{k}=i\sin(kN)$.
The quasiparticle energy is $\mathcal{E}_{k}=|\sin(kN)|$, and $|u_{k}|=\frac{1}{\sqrt{2}},|v_{k}|=\frac{1}{\sqrt{2}}$. The crucial ratio is determined by
$\frac{v_{k}}{u_{k}}=\frac{\Delta_{k}}{0+E_{k}}=\frac{i\sin(kN)}{|\sin(kN)|}=i\cdot\text{sgn}(\sin(kN))\,.$ (S124)

### v2

@@ D.3 The BCS representation of the crosscap state
$H_{\text{parent}}=-\sum_{k>0}\Psi_{k}^{\dagger}\mathbf{H}_{k}\Psi_{k},$ (S118)
@@
$\mathbf{H}_{k}=\begin{pmatrix}\epsilon_{k}^{0}&\Delta_{k}^{0}\\ \Delta_{k}^{0*}&-\epsilon_{k}\end{pmatrix}=\begin{pmatrix}0&i\sin(kN)\\ -i\sin(kN)&0\end{pmatrix}\,$. (S119)
Here, we’ve used the specific values for our parent Hamiltonian: the diagonal ”energy” term is $\epsilon_{k}^{0}=0$, and the off-diagonal ”pairing” term is $\Delta_{k}=i\sin(kN)$.
@@
We find the coefficients by substituting the transformation into the Hamiltonian and demanding that all off-diagonal terms like $b_{k}^{\dagger}b_{-k}^{\dagger}$ and $b_{k}b_{-k}$ vanish. This leads to a standard set of equations that relate the coefficients to the Hamiltonian parameters $\epsilon_{k}^{0}$ and $\Delta_{k}^{0}$.
@@
$\mathcal{E}_{k}^{0}=\sqrt{(\epsilon_{k}^{0})^{2}+|\Delta_{k}^{0}|^{2}}$ (S121)
@@
$|u_{k}|^{2}=\frac{1}{2}\left(1+\frac{\epsilon_{k}^{0}}{\mathcal{E}_{k}^{0}}\right)\quad\text{and}\quad|v_{k}|^{2}=\frac{1}{2}\left(1-\frac{\epsilon_{k}^{0}}{\mathcal{E}_{k}^{0}}\right)\,$, (S122)
@@
$\frac{v_{k}}{u_{k}}=\frac{\Delta_{k}^{0}}{\epsilon_{k}^{0}+\mathcal{E}_{k}^{0}}\,.$ (S123)
Now we substitute the specific parameters for the crosscap state’s parent Hamiltonian: $\epsilon_{k}^{0}=0$ and $\Delta_{k}^{0}=i\sin(kN)$.
The quasiparticle energy is $\mathcal{E}_{k}^{0}=|\sin(kN)|$, and $|u_{k}|=\frac{1}{\sqrt{2}},|v_{k}|=\frac{1}{\sqrt{2}}$. The crucial ratio is determined by
$\frac{v_{k}}{u_{k}}=\frac{\Delta_{k}^{0}}{0+\mathcal{E}_{k}^{0}}=\frac{i\sin(kN)}{|\sin(kN)|}=i\cdot\text{sgn}(\sin(kN))\,.$ (S124)
@@ D.4 The Fermi-Dirac distribution from TPQ states
D.5 The parent Hamiltonian for the TPQ State
We seek a parent Hamiltonian, $H_{\text{parent}}$, whose ground state is $|\Psi_{\beta}\rangle$. Since $|\Psi_{\beta}\rangle$ is a Gaussian state, the most general quadratic parent Hamiltonian can be written in the Nambu spinor basis $\Psi_{k}^{\dagger}=\begin{pmatrix}c_{k}^{\dagger}&c_{-k}\end{pmatrix}$ as:
$H_{\text{parent}}(\beta)=\sum_{k}\Psi_{k}^{\dagger}H_{k}\Psi_{k}=\sum_{k}\begin{pmatrix}c_{k}^{\dagger}&c_{-k}\end{pmatrix}\begin{pmatrix}\epsilon_{k}&\Delta_{k}\\ \Delta^{*}_{k}&-\epsilon_{k}\end{pmatrix}\begin{pmatrix}c_{k}\\ c_{-k}^{\dagger}\end{pmatrix}$. (S139)
This expands to the form
$H_{\text{parent}}(\beta)=\sum_{k}\epsilon_{k}(c_{k}^{\dagger}c_{k}-c_{-k}c_{-k}^{\dagger})+\sum_{k}\left(\Delta_{k}c_{k}^{\dagger}c_{-k}^{\dagger}+\Delta^{*}_{k}c_{-k}c_{k}\right),$ (S140)
where $\epsilon_{k}$ is the single-particle energy spectrum and $\Delta_{k}$ is the pairing potential of the parent Hamiltonian. Our goal is to find the functions $\xi_{k}$ and $\Delta_{k}$ that produce $|\Psi_{\beta}\rangle$ as the ground state.
For a general BCS Hamiltonian of the form in Eq. (S139), the occupation probability of its ground state is given by the standard relation:
$|v_{k}|^{2}=\frac{1}{2}\left(1-\frac{\epsilon_{k}}{\mathcal{E}_{k}}\right),$ (S141)
where $\mathcal{E}_{k}=\sqrt{(\epsilon_{k})^{2}+|\Delta_{k}|^{2}}$ is the quasiparticle energy spectrum of the parent Hamiltonian. We now enforce that this ground state is our TPQ state by equating Eq. (S141) with Eq. (S137), which allows us to solve for the ratio $\epsilon_{k}/\mathcal{E}_{k}$ as
$\frac{\epsilon_{k}}{\mathcal{E}_{k}}=\tanh\left(\frac{\beta E(k)}{2}\right).$ (S142)
This equation provides a fundamental constraint on the parameters of our desired parent Hamiltonian. To solve for $\epsilon_{k}$ and $\Delta_{k}$ separately, we make the natural choice of setting the single-particle spectrum of the parent Hamiltonian equal to that of the physical XX chain, i.e., $\epsilon_{k}=E(k)$. With this assumption, the quasiparticle energy of the parent Hamiltonian is determined from Eq. (S142):
$\mathcal{E}_{k}=\frac{\epsilon_{k}}{\tanh\left(\frac{\beta E(k)}{2}\right)}=E(k)\coth\left(\frac{\beta E(k)}{2}\right).$ (S143)
Finally, we use the definition $\mathcal{E}_{k}^{2}=(\epsilon_{k})^{2}+|\Delta_{k}|^{2}$ to find the required pairing potential:
$|\Delta_{k}|^{2}$ $=\mathcal{E}_{k}^{2}-(\epsilon_{k})^{2}$
$=E(k)^{2}\coth^{2}\left(\frac{\beta E(k)}{2}\right)-E(k)^{2}$
$=E(k)^{2}\left[\coth^{2}\left(\frac{\beta E(k)}{2}\right)-1\right]=\frac{E(k)^{2}}{\sinh^{2}\left(\frac{\beta E(k)}{2}\right)}.$ (S144)
This gives the magnitude of the pairing potential as $|\Delta_{k}|=|E(k)|/\left|\sinh\left(\frac{\beta E(k)}{2}\right)\right|$.